Suppose a family's annual income is $58,000; if the marginal propensity to save (MPS) is 0.25, and the income for the family decreases by $15,000, then the decrease in consumption will be
Select one:
a. $3,750.
b. $10,500.
c. $11,250.
d. $1,500
1
Expert's answer
2016-02-11T00:00:47-0500
I1 = $58,000; MPS = 0.25, I2 - I1 = $15,000.
MPC = 1 - MPS = 0.75,
MPC = (C2 - C1)/(I2 - I1),
The decrease in consumption will be (C2 - C1) = MPC*(I2 - I1) = 0.75*15,000 = $11,250.
So, the right answer is c. $11,250.